OpenStudy (anonymous):

simplify. Assume that no denominator is equal to zero. 7^13\7^3

OpenStudy (cj7529):

use the general rule in indicies that X^A / X^B = X^(A-B) So, 7^13/7^3 = 7^(13-3) =7^10

OpenStudy (anonymous):

thank you!!!